what is the sum in unit vector notation of the two vector a=5i^+3j^ and b=-3i^+2j^? what is the magnitude and diraction of a+b?
 
 
 
 

Grade:12th pass

1 Answers

Khimraj
5 years ago
a = 5\hat{i} + 3\hat{j}
b = -3\hat{i} + 2\hat{j}
a + b = 2\hat{i} + 5\hat{j}
magnitude = \sqrt{2^{2}+5^{2}} = \sqrt{29}
direction = (2\hat{i} + 5\hat{j})/\sqrt{29}
Hope it clears..............................